OpenStudy (anonymous):

Find the exact value of each logarithm log10 √10 can someone show me the steps

OpenStudy (anonymous):

\[log_x(x^{1/3}) = \frac{1}{3}*log_x{x}\] \[log_x{x} = 1\]

OpenStudy (anonymous):

\[\log10\sqrt{10}\] or \[\log _{10} 10\sqrt{10}\]

OpenStudy (anonymous):

it wouldnt really matter anyways

OpenStudy (anonymous):

it is \[log_{10}(10)\]

OpenStudy (anonymous):

it should be sqrt(10) in the parentheses. forgot to type that.

OpenStudy (anonymous):

my first post should get you the answer very easily.